# Forecasting the U.S. Renewable-Energy Mix with an ALR-BDARMA Compositional Time-Series Framework

### v3

@@ Abstract
Accurate forecasts of the U.S. renewable-generation mix are essential for planning transmission upgrades, sizing storage, and setting balancing-market rules. We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025. The mean vector is modeled with a parsimonious VAR(2) in additive-log-ratio space, while the Dirichlet concentration parameter follows an intercept plus ten Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion.
Forecast performance is assessed with a 61-split rolling-origin experiment that issues twelve-month density forecasts from January 2019 to January 2024. Compared with three alternatives—a Gaussian VAR(2) fitted in transform space, a seasonal naïve that repeats last year’s proportions, and a drift-free ALR random walk—BDARMA lowers the mean continuous-ranked probability score by 15–60 %, achieves component-wise 90 % interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward. These results highlight BDARMA’s ability to deliver sharp, well-calibrated probabilistic forecasts for multivariate renewable-energy shares without sacrificing point precision.
Keywords: compositional time series; Dirichlet state-space; Bayesian forecasting; renewable-energy mix; seasonality.
@@ 1 Introduction
Electric‑sector decarbonization hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve. Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S. Energy Information Administration, 2024; U.S. Energy Information Administration, 2025). Renewables already supply about one‑fifth of U.S. utility‑scale electricity and their share is expected to double before 2050, making the coherence and accuracy of share forecasts more important than ever.
Why compositional methods?
@@
Simplex‑native likelihoods.
Recent research therefore models the composition itself. A Dirichlet ARMA process was proposed by Zheng et al. (2017), a dynamic Dirichlet–multinomial filter by Koopman et al. (2023), and a deep hierarchical Dirichlet forecaster by Das et al. (2023). In the cross‑sectional domain, Morais et al. (2018) show that Dirichlet and compositional regression can outperform traditional attraction models when explaining brand market shares, underscoring the versatility of simplex‑based methods. A direct antecedent to the present study is the Bayesian Dirichlet ARMA (BDARMA) framework introduced by Katz et al. (2024), and subsequently explored with shrinkage priors for trading‑sector shares by Katz et al. (2025), which we adopt here as the data model for a new application to the U.S. renewable‑energy mix.
Evidence from energy studies.
@@
Contribution.
We apply the Bayesian Dirichlet ARMA framework to the seven‑component U.S. renewable‑energy mix measured monthly from 2010 to 2024, a data set with pronounced seasonality and secular trends hitherto unaddressed in the Dirichlet literature. Forecast skill is benchmarked against three alternatives: a Gaussian VAR(2) in additive‑log‑ratio space with identical Fourier dummies, a seasonal naïve that repeats the mix observed twelve months earlier, and a drift‑free ALR random walk. A 61‑split rolling protocol produces 732 out‑of‑sample density forecasts and shows that the Dirichlet model attains the lowest continuous‑ranked probability score at every horizon while matching the VAR’s point accuracy. A full‑sample forecast issued in January 2025 encloses a trajectory where wind and solar exceed one‑third of renewable generation by early 2026, with credible bands roughly half as wide as those implied by the VAR—information directly relevant for transmission and storage planning.
The remainder of the paper is organized as follows. Section 2 describes the EIA data and seasonal covariates. Section 3 presents the BDARMA and benchmark models. Section 4 details the rolling evaluation protocol and scoring rules. Results are discussed in Section 5, and Section 6 concludes with policy implications and avenues for future research.
@@ 2.1 Exploratory data analysis
Cross‑source dependence.
Figure 3 (a) plots the correlation matrix of the six ALR coordinates $e_{1},\dots,e_{6}$ (biofuels as reference). Solar and wind move almost one‑for‑one relative to biofuels ($\rho_{e_{5},e_{6}}=0.97$), whereas hydro and wind are strongly anti‑correlated ($\rho\approx-0.86$). Panel (b) confirms these pair‑wise relations are non‑linear, displaying the characteristic banana‑shaped clouds induced by log‑ratio geometry.
Distributional heterogeneity.
@@
(a) Seasonal box plots by component
(b) Average intra‑year profile
@@
(a) Correlation matrix of ALR coordinates
(b) Pairwise ALR scatter plots
@@
Figure 5: ACF of squared residuals for the preferred VAR(2) + season model. Modest spikes in $e_{2}$ and $e_{5}$ motivate the component‑wise GARCH precision adopted later.
Because the data exhibit markedly different seasonal amplitudes, strong yet uneven cross‑correlations, and component‑specific dispersion and because the residual diagnostics show that two lags are the minimum needed for whiteness, we adopt a specification built on three mutually reinforcing elements: a second‑order vector autoregressive mean in ALR space to capture short‑run dynamics, a single seasonal precision curve that modulates uncertainty across the calendar year for every component, and a Dirichlet likelihood that both imposes the compositional sum‑to‑one constraint and allows each source’s variance to remain distinct.
Software environment.
All computations were carried out in R 4.3.2 (R Core Team, 2024) with Stan 2.33 via the rstan interface (Goodrich et al., 2024; Stan Development Team, 2023). Data wrangling, graphics and tables relied on tidyverse (Wickham et al., 2019), lubridate (Grolemund and Wickham, 2011), janitor (Firke, 2023), scales (Wickham and Seidel, 2019), patchwork (Pedersen, 2025), ggcorrplot (Kassambara, 2022), GGally (Di Cook et al., 2021) and kableExtra (Zhu, 2021). Compositional methods used compositions (van den Boogaart and Tolosana‐Delgado, 2024) and the transport package for Aitchison norms (Schuhmacher et al., 2020). Time‑series estimation and testing employed vars (Pfaff, 2008), FinTS (Pfaff, 2024) and MTS (Tsay, 2023).
@@ 3 Forecasting model
Seasonal precision curve.
Calendar variation in forecast dispersion is captured by letting the log‑precision depend on an intercept and ten Fourier harmonics:
$\log\varphi_{t}=\mathbf{f}_{t}^{\!\top}\boldsymbol{\gamma},\quad\mathbf{f}_{t}=(1,\mathbf{g}_{t}^{\!\top})^{\!\top},\;\mathbf{g}_{t}=\bigl(\sin\tfrac{2\pi t}{12},\cos\tfrac{2\pi t}{12},\ldots,\sin\tfrac{10\pi t}{12},\cos\tfrac{10\pi t}{12}\bigr)^{\!\top},\quad\boldsymbol{\gamma}\in\mathbb{R}^{11}.$ (3)
State (mean) dynamics.
@@ 4.1 Generating predictive distributions
Let $\widehat{\boldsymbol{\eta}}_{s,h}$ and $\widehat{\mathbf{V}}_{s,h}$ be, respectively, the conditional mean and covariance of the Gaussian forecast for the ALR vector at horizon $h$. We draw $\boldsymbol{\eta}^{(m)}_{s,h}\sim\mathcal{N}(\widehat{\boldsymbol{\eta}}_{s,h},\widehat{\mathbf{V}}_{s,h})$, transform with $\text{alr}^{-1}$, and obtain $\mathcal{P}^{\text{tVAR}}_{s,h}=\{\text{alr}^{-1}(\boldsymbol{\eta}^{(m)}_{s,h})\}_{m=1}^{M}.$ Because the back‑transformation preserves the unit sum.
@@ 4.2 Scoring rules
Continuous‑ranked probability score.
Writing $\|\mathbf{a}\|_{1}=\sum_{j=1}^{7}|a_{j}|$ for the $\ell_{1}$ norm, the sample‑based estimator is
$\operatorname{CRPS}_{s,h}(\mathcal{P})=\frac{1}{M}\sum_{m=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{s,h}\bigr\|_{1}-\frac{1}{2M^{2}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{\,s,h}^{(m^{\prime})}\bigr\|_{1},$ (5)
where $\mathcal{P}$ denotes the empirical draw set for the model in question. The CRPS integrates the Brier score over all threshold events inside the simplex and is measured in share points.
@@ 5.1 Forecast accuracy across horizons
Table 3 reports the mean continuous‑ranked probability score (CRPS) and Table 4 the mean Aitchison root‑mean‑square error (RMSE) for every horizon; the same information is visualised in Figures 6 and 7. When scores are aggregated over all sixty‑one rolling origins, the Bayesian‑Dirichlet ARMA (BDARMA) dominates the field. Its mean CRPS at one step ahead is  0.0045\,0.0045, which is roughly one quarter lower than the transform‑space VAR(2) (tVAR(2)) and more than fifty percent lower than either naïve rule. The advantage widens as the horizon lengthens; by month 12, BDARMA still improves on tVAR(2) by around twenty percent and on the seasonal naïve (S‑NAIVE) by a factor of 1.4, while the additive‑log‑ratio random walk (ALR‑RW) remains the weakest performer throughout.
Point‑forecast results tell a complementary story. Up to eight months ahead, BDARMA and tVAR(2) deliver virtually identical RMSEs (Figure 7); thereafter the Gaussian VAR displays a slight numerical edge that peaks at just one‑hundredth of an Aitchison unit. That edge, however, is achieved at the cost of substantially wider predictive intervals, as reflected in the persistently higher CRPS. Both naïve rules fare considerably worse: S‑NAIVE holds a constant RMSE of about $0.15$ and a CRPS that is almost double BDARMA’s, whereas ALR‑RW deteriorates sharply until mid‑year before converging towards the other benchmarks only because its naive dispersion collapses.
Taken together, these findings underline three practical lessons. First, seasonality matters: the harmonic regressors embedded in $\mathbf{f}_{t}$ provide an immediate and durable boost over models that ignore calendar structure. Second, coherence alone is not enough; although tVAR(2) respects the simplex, its Gaussian scale inflates forecast spread and leaves roughly one‑fifth of the potential CRPS improvement on the table. Finally, the Dirichlet precision curve in BDARMA succeeds in producing densities that are simultaneously sharp, well calibrated and competitive in point terms even at the longest horizon considered.
@@
Figure 7: Mean Aitchison RMSE by horizon; lower values indicate more accurate point forecasts.
@@ 5.2 Coverage of BDARMA predictive intervals
The Monte‑Carlo intervals are well calibrated. Table 5 records the proportion of split‑horizon combinations where the realised share vector falls inside the component‑wise 90% band. Coverage rises from 86% at one month to 99% at a full year; the geometric ceiling for joint inclusion is 48% at $h=1$, so the upward trend is expected. Component averages range from 86% for biofuels to 100% for geothermal (Table 6), giving an overall mean of 94.8%.
Table 5: Empirical 90 % coverage of BDARMA component intervals.
@@
Table 6: Component‑wise BDARMA coverage ($61\times 12$ forecasts).
@@ 5.3 Fixed‑origin comparison
Figure 8: Twelve‑month forecasts issued 2025‑01‑01 after refitting all models to the 2010–2024 sample. Blue shading denotes the BDARMA 90 % predictive interval; colored lines are posterior or plug‑in medians. Axes are free by facet.
Figure 8 shows the one‑year‐ahead trajectories for each technology by model. The central messages are straightforward.
First, BDARMA delivers not only a median path but also a well calibrated 90 % interval. For planners this means that the same forecast set supports both point estimates—needed, for instance, in generation‑mix optimization—and a quantification of downside or upside risk that is essential when determining reserve margins, storage dispatch protocols or hedging requirements. The competing methods provide no such density information, forcing practitioners to bolt on ad‑hoc error bands or run scenario analyses that may or may not respect the unit‑sum constraint.
Second, because BDARMA’s concentration parameter is allowed to vary seasonally, the width of its interval changes appreciably across components. High‑growth and strongly seasonal sources such as wind and solar retain non‑negligible uncertainty, implying that transmission and balancing services must still accommodate significant swings even as their average shares rise. In contrast, geothermal and waste show near‑deterministic bands, suggesting that policy targets focused on these technologies can be stated with greater confidence and narrower tolerances.
Third, where the medians diverge the direction is informative in its own right. The harmonic naïve inherits last year’s level and thereby under‑projects solar in the presence of an upward trend; the random walk does the opposite by ignoring seasonality entirely; and the Gaussian VAR, although often close to BDARMA in the center, cannot reveal whether any gap that remains is material relative to forecast dispersion. From a policy standpoint this matters when setting technology‑specific incentives. Under‑estimating solar growth, for example, risks under‑sizing flexible resources needed to absorb midday surpluses, whereas over‑estimating hydro stability can mask the need for additional spring reserves.

### v4

@@ Abstract
Accurate forecasts of the U.S. renewable energy consumption mix are essential for planning transmission upgrades, sizing storage, and setting balancing market rules. We introduce a Bayesian Dirichlet ARMA model (BDARMA) tailored to monthly shares of hydro, geothermal, solar, wind, wood, municipal waste, and biofuels from January 2010 through January 2025. The mean vector is modeled with a parsimonious VAR(2) in additive log ratio space, while the Dirichlet concentration parameter follows an intercept plus five Fourier harmonics, allowing seasonal widening and narrowing of predictive dispersion. Forecast performance is assessed with a 61 split rolling origin experiment that issues twelve month density forecasts from January 2019 to January 2024. Compared with three alternatives, a Gaussian VAR(2) fitted in transform space, a seasonal naive that repeats last year’s proportions, and a drift free ALR random walk, BDARMA lowers the mean continuous ranked probability score by 15 to 60 percent, achieves component wise 90 percent interval coverage near nominal, and maintains point accuracy (Aitchison RMSE) on par with the Gaussian VAR through eight months and within 0.02 units afterward. These results highlight BDARMA’s ability to deliver sharp, well calibrated probabilistic forecasts for multivariate renewable energy shares without sacrificing point precision. Supplementary Material: An electric power only robustness analysis aligned with the EIA STEO baseline (strict vintaging) is reported in Sections S1 and S2.
Keywords: Compositional time series; Dirichlet state-space; Bayesian forecasting; renewable energy mix; seasonality.
@@ 1 Introduction
Electric‑sector decarbonization hinges not only on expanding renewable output but also on anticipating how the mix of generation technologies will evolve. Hydropower, wind, solar, biomass and geothermal differ sharply in marginal cost, intermittency and siting constraints; reliable medium‑term mix forecasts therefore shape transmission expansion, storage sizing and market design (International Energy Agency, 2024; U.S. Energy Information Administration, 2024, 2025). Renewables already supply about one‑fifth of U.S. utility‑scale electricity and their share is expected to double before 2050, making the coherence and accuracy of share forecasts more important than ever.
@@
Recent research therefore models the composition itself. A Dirichlet ARMA process was proposed by Zheng et al. (2017), a dynamic Dirichlet–multinomial filter by Koopman et al. (2023), and a deep hierarchical Dirichlet forecaster by Das et al. (2023). In the cross‑sectional domain, Morais et al. (2018) show that Dirichlet and compositional regression can outperform traditional attraction models when explaining brand market shares, underscoring the versatility of simplex‑based methods. A direct antecedent to the present study is the Bayesian Dirichlet ARMA (BDARMA) framework (Katz et al., 2024; Katz and Weiss, 2025), and subsequently explored with shrinkage priors for trading‑sector shares by Katz et al. (2025), which we adopt here as the data model for a new application to the U.S. renewable‑energy mix.
@@
We apply the Bayesian Dirichlet ARMA framework to the seven‑component U.S. renewable‑energy mix measured monthly from 2010 to 2024, a data set with pronounced seasonality and secular trends hitherto unaddressed in the Dirichlet literature. Forecast skill is benchmarked against three alternatives: a Gaussian VAR(2) in additive‑log‑ratio space with identical Fourier dummies, a seasonal naïve that repeats the mix observed twelve months earlier, and a drift‑free ALR random walk. A 61‑split rolling protocol produces 732 out‑of‑sample density forecasts and shows that the Dirichlet model attains the strongest probabilistic performance while maintaining the VAR’s point accuracy. A full-sample forecast to early 2026 projects wind and solar surpassing one-third of renewable generation, providing a coherent picture for transmission and storage planning.
The remainder of the paper is organized as follows. Section 2 describes the EIA data and seasonal covariates. Section 3 presents the BDARMA and benchmark models. Section 4 details the rolling evaluation protocol and scoring rules. Results are discussed in Section 5, and Section 6 concludes with policy implications and avenues for future research. A complementary robustness analysis aligned with the EIA STEO definitions is presented in the Supplementary Material (Secs. S1–S2).
@@ 2 Data
Electric‑power–only benchmarking.
For comparisons to the EIASTEO industry baseline we also construct an electric‑power–only view (hydro, geothermal, solar = utility‑scale + small PV, wind, wood, waste) with monthly closure to the simplex. Implementation details and the strict vintaging rule are in the Supplementary Material (Sec. S1).
@@ 2.1 Exploratory data analysis
Figure 3 (a) plots the correlation matrix of the six ALR coordinates $e_{1},\dots,e_{6}$ (biofuels as reference). Solar and wind move almost one‑for‑one relative to biofuels ($\rho_{e_{3},e_{4}}=0.97$), whereas hydro and wind are strongly anti‑correlated ($\rho_{e_{1},e_{4}}\approx-0.86$). Panel (b) confirms these pair‑wise relations are non‑linear, displaying the characteristic banana‑shaped clouds induced by log‑ratio geometry.
@@
Figure 5: ACF of squared residuals for the preferred VAR(2) + season model.
[…]
All computations were carried out in R 4.3.2 with Stan 2.33 via the cmdstanr interface. (Goodrich et al., 2024; Stan Development Team, 2023). Data wrangling, graphics and tables relied on tidyverse (Wickham et al., 2019), lubridate (Grolemund and Wickham, 2011), janitor (Firke, 2023), scales (Wickham and Seidel, 2019), patchwork (Pedersen, 2025), ggcorrplot (Kassambara, 2022), GGally (Di Cook et al., 2021) and kableExtra (Zhu, 2021). Compositional methods used compositions (van den Boogaart and Tolosana‐Delgado, 2024) and the transport package for Aitchison norms (Schuhmacher et al., 2020). Time‑series estimation and testing employed vars (Pfaff, 2008), FinTS (Pfaff, 2024) and MTS (Tsay, 2023).
@@ 3 Forecasting model
Calendar variation in forecast dispersion is captured by letting the log‑precision depend on an intercept and five Fourier harmonics (ten sine/cosine terms):
$\log\phi_{t}=\mathbf{f}_{t}^{\!\top}\boldsymbol{\gamma},\quad\mathbf{f}_{t}=(1,\mathbf{g}_{t}^{\!\top})^{\!\top},\;\mathbf{g}_{t}=\bigl(\sin\tfrac{2\pi t}{12},\cos\tfrac{2\pi t}{12},\ldots,\sin\tfrac{10\pi t}{12},\cos\tfrac{10\pi t}{12}\bigr)^{\!\top},\quad\boldsymbol{\gamma}\in\mathbb{R}^{11}.$ (3)
@@
With a scalar precision $\phi_{t}$, the Dirichlet implies a restricted covariance: $\mathrm{Cov}(y_{i},y_{j})=-\mu_{i}\mu_{j}/(\phi_{t}+1)$ for $i\neq j$, i.e., negative off‑diagonals of fixed shape. Cross‑component comovement beyond the unit‑sum constraint therefore enters through the mean dynamics rather than the observation variance. Extensions include generalized Dirichlet or logistic‑normal layers, or component‑specific precisions $\phi_{j,t}$ with regularization; we leave these for future work.
@@ 4.1 Generating predictive distributions
Let $\widehat{\boldsymbol{\eta}}_{s,h}$ and $\widehat{\mathbf{V}}_{s,h}$ be, respectively, the conditional mean and covariance of the Gaussian forecast for the ALR vector at horizon $h$. We draw $\boldsymbol{\eta}^{(m)}_{s,h}\sim\mathcal{N}(\widehat{\boldsymbol{\eta}}_{s,h},\widehat{\mathbf{V}}_{s,h})$, transform with $\text{alr}^{-1}$, and obtain $\mathcal{P}^{\text{tVAR}}_{s,h}=\{\text{alr}^{-1}(\boldsymbol{\eta}^{(m)}_{s,h})\}_{m=1}^{M}.$ Multi‑step forecasts are generated under the Gaussian‑innovation assumption and then mapped back to shares with $\operatorname{alr}^{-1}$, which preserves unit‑sum coherence.
@@ 4.2 Scoring rules
Energy score (multivariate CRPS).
Writing $\|\mathbf{a}\|_{1}=\sum_{j=1}^{7}|a_{j}|$ for the $\ell_{1}$ norm, the sample‑based energy score (ES; a multivariate generalization of the CRPS) is
$\operatorname{ES}_{s,h}(\mathcal{P})=\frac{1}{M}\sum_{m=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{s,h}\bigr\|_{1}-\frac{1}{2M^{2}}\sum_{m=1}^{M}\sum_{m^{\prime}=1}^{M}\bigl\|\mathbf{y}_{\,s,h}^{(m)}-\mathbf{y}_{\,s,h}^{(m^{\prime})}\bigr\|_{1}.$ (5)
We use the $\ell_{1}$ norm so that units are “share points”; ES remains a strictly proper scoring rule under common norms.
@@
4.3 External baseline and scoring in electric‑only space
A matched evaluation against the vintaged EIASTEO baseline in the EP‑only frame uses the same rolling origins, horizons, and scoring rules; see Supplementary Sec. S1 for construction and Sec. S2 for the horizon‑by‑horizon results (Figs. S1–S2; Tables S1–S2).
@@ 5.1 Forecast accuracy across horizons
Tables 3 and 4 report mean CRPS and mean Aitchison RMSE by horizon; Figures 6 and 7 visualize the same quantities. Across sixty‑one rolling origins, the Bayesian Dirichlet ARMA (BDARMA) model attains the lowest CRPS at every horizon. At one month it is about one quarter lower than the transform‑space VAR(2) and more than half lower than either naïve rule, and the advantage widens with lead: by twelve months BDARMA still improves on tVAR(2) by roughly one fifth and on S‑NAIVE by about forty percent, while ALR‑RW remains weakest overall.
Point errors give a complementary view. Through eight months, BDARMA and tVAR(2) yield nearly identical RMSEs; from month nine onward, the Gaussian VAR gains a small edge that peaks at roughly one hundredth of an Aitchison unit. That edge comes with broader predictive spreads, which shows up as persistently higher CRPS for the VAR.
Technology‑specific interpretation.
Component patterns matter for planning. Wind and hydro exhibit the largest seasonal swings, so their predictive bands are wider and more seasonally structured, informing spring runoff scheduling for hydro and winter ramping reserves for wind. Solar’s long‑run rise with a summer plateau produces medium‑horizon gains that help quantify midday surplus risk and storage sizing. Geothermal and waste behave almost deterministically, supporting narrow tolerance bands for compliance or procurement. Wood remains comparatively volatile across horizons, arguing for conservative hedging where biomass supply or policy constraints bind. These qualitative statements align with Table 6, where geothermal and waste approach perfect inclusion while wind and biofuels are harder to capture due to stronger seasonality and policy or demand variability.
@@
Figure 7: Mean Aitchison RMSE by horizon. Lower values indicate more accurate point forecasts.
@@ 5.2 Coverage of BDARMA predictive intervals
The Monte Carlo intervals are well calibrated. Componentwise 90% coverage rises from 86% at one month to 99% by a full year (Table 5). By technology (Table 6), geothermal and waste approach perfect inclusion; solar and hydro are very close to nominal; wind and biofuels are lower, consistent with larger seasonal amplitude and policy‑ or demand‑driven swings.
Table 5: Empirical 90 percent coverage of BDARMA component intervals.
@@
Table 6: Componentwise BDARMA coverage (61 by 12 forecasts).
@@ 5.3 Fixed‑origin comparison
Figure 8: Twelve‑month forecasts issued 2025-01-01 after refitting all models to the 2010–2024 sample. Blue shading denotes the BDARMA 90% predictive interval. Colored lines are posterior or plug‑in medians. Axes are free by facet.
The one‑year trajectories in Figure 8 reinforce these patterns. BDARMA produces calibrated bands that widen where seasonality and trend are strong and tighten where series are stable, supporting a single forecast set for both point planning and risk assessment. The naïve rules supply medians only. The Gaussian VAR yields medians close to BDARMA but cannot indicate whether any remaining gap is material relative to forecast dispersion.
If decisions depend almost entirely on point forecasts, exogenous regressors are abundant and trusted, and speed is critical, the transform‑space VAR(2) is reasonable, especially beyond nine months where it has a slight RMSE edge. The seasonal naïve can suffice for steady components and as a monitoring benchmark where deviations from a baseline are the main signal. BDARMA remains preferable when calibrated densities, compositional coherence, and seasonal uncertainty are central to the decision.
Sequence models such as LSTMs or GRUs can be competitive with long histories and rich covariates. In monthly, medium‑length settings with few components, they require careful constraints to preserve the unit sum and substantial tuning to obtain calibrated densities. Relative to such sequence learners and to gradient‑boosted ensembles, BDARMA offers three practical advantages for this task: (i) coherence by construction (no post‑hoc renormalization), (ii) direct density forecasts in share space rather than ad‑hoc bands, and (iii) strong performance with limited history. The approaches are complementary: neural or boosted summaries of weather or policy can enter BDARMA as exogenous features, and neural forecasts can be used as external signals.
Estimation uses four chains with 1000 warm‑up and draw counts; origins and chains parallelize across cores. Design matrices grow linearly with the number of components and harmonics.
A matched electric‑power‑only comparison against the vintaged EIASTEO baseline confirms the main findings: BDARMA leads at one month, while STEO is strongest beyond two months. Details are in the Supplementary Material (Sections S1–S2; Figures S1–S2; Tables S1–S2).